Solve the linear equations:
x + 3y + 2z = -3
x + y + z = -1
3x + 5y + 6z = -5

Answer by richwmiller(17219)

x + 3y + 2z = -3
x + y + z = -1
3x + 5y + 6z = -5
x + 3y + 2z = -3
3x + 9y + 6z = -9
3x + 5y + 6z = -5
4y=-4
y=-1
x + 3y + 2z = -3
x + y + z = -1
2y+z=-2
-2+z=-2
z=0
x + y + z = -1
x-1+0=-1
x=0
x=0
y=-1
z=0
(0,-1,0)
Warning: I am still sick in bed. Double check my work.